OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) –4 16 0 0 4 –16 8 –32

OpenStudy (jlastino):

Ok I'll rewrite that as ordered pairs (x,f(x)) (-4,16) (0,0) (4,-16) (8,-32) Just like before, use the two point formula Pick any two points from the 4 ( I'll take (-4,16) and (0,0) so x1= -4 x2= 0 y1=16 y2=0 y-y1= ((y2-y1)/(x2-x1)) (x-x1) y-16= (0-16)/(0+4) (x+4) y-16= -16/+4 (x+4) y-16= -4 (x+4) y-16= -4x-16 y=-4x f(x)=-4x To check if this is true try x= -4 -4(-4)= 16 ---> value of f(x)

OpenStudy (anonymous):

thankyou